# The influence of cell phenotype on collective cell invasion into the extracellular matrix

### v1

@@ 4.2 Non-ECM-modulating phenotypes display restricted invasion into dense collagen environments
Another important prediction of the model is that the initial collagen fibre distribution indirectly influences cell proliferation. This occurs because the cell clusters are more compact when the collagen fibres are more dense and more aligned (Figure 3(c),(d), and Figure S2 in the Supplementary Information). This effect is not apparent until approximately $24$ hours of (Figure 3(e),(f)), which corresponds to the average cell cycle length.
@@ 4.3 Collagen-modulating phenotypes display different invasion patterns
We now examine the dynamics of a collagen-secreting phenotype, setting the degradation rate to be zero, $d=0.0$, and the initial total area fraction of collagen to be low so that there is substantial area available into which cells can secrete new collagen. Here, the remodelling process is more intricate, and displays spatial variations across the cluster (see Figure 4(b)). For cells in the top and bottom regions (as defined in Figure 2), migration is primarily aligned with the vertically oriented major fibre direction. In these regions, anisotropy increases over time as newly secreted fibres reinforce this vertical alignment. For cells in the left and right regions (as defined in Figure 2), migration is primarily driven by population pressure arising from cell–cell interactions, and is aligned with the horizontally oriented minor fibre direction. As the cells secrete new collagen, which is primarily horizontally aligned, the fibre anisotropy initially decreases over time as the fibre distribution becomes more uniform (i.e., $\lambda_{1}\approx\lambda_{2}$). It then increases as newly secreted fibres are continually laid down in the horizontal alignment, so that the major fibre direction becomes horizontal. In the central region, the anisotropy decreases over time as cells move predominantly randomly.
@@ 4.4 Tensioning secretion and degradation in the phenotype alters invasion patterns
To investigate how a phenotype that both secretes and degrades fibres affects collagen fibre dynamics and cell invasiveness, we perform a parameter sweep across different fibre secretion and degradation rates. We consider a scenario in which there is very strong contact guidance initially due to a high area fraction of aligned collagen. Figure 5(a) illustrates the cell and collagen fibre distributions for small secretion with large degradation rates ($\star^{1}$), large secretion and large degradation rates ($\star^{2}$), and large secretion with small degradation rates ($\star^{3}$). For a cell phenotype that predominantly degrades collagen ($\star^{1}$), the initial collagen distribution is degraded, reducing the strength of contact guidance while facilitating cell invasion in all directions. This is consistent with results in Figure 4(a). For a cell phenotype that equally secretes and degrades collagen ($\star^{2}$), the initially aligned fibres are degraded: this results in reduced contact guidance so that newly secreted collagen is relatively isotropic. Finally, for a cell phenotype that predominantly secretes collagen ($\star^{3}$), the initial anisotropy, and thus the contact guidance cues, are reinforced in the top and bottom regions. This because cells in these regions tend to migrate vertically due to population pressure, even in the absence of contact guidance. This promotes enhanced invasion along the vertical axis, while horizontal invasion is suppressed because guidance cues from population pressure are in tension with those from the collagen distribution. This is consistent with results in Figure 4(b).
@@
Figure 5: Tensioning secretion and degradation in the phenotype alters invasion patterns. Results from a parameter sweep across secretion ($s$) and degradation ($d$) rates. (a) Cell and collagen fibre distributions at $42$ hours with $\star^{1}$$\left(s,d\right)=\left(2^{-10},2^{-4}\right)$ $\rm{min}^{-1}$, $\star^{2}$$\left(s,d\right)=\left(2^{-4},2^{-4}\right)$ $\rm{min}^{-1}$, and $\star^{3}$$\left(s,d\right)=\left(2^{-4},2^{-10}\right)$ $\rm{min}^{-1}$, respectively. (b) Heat map showing the similarity of the average cell density along the horizontal direction, compared to that of the non-degrading cell phenotype ($d=0,\rm{min}^{-1}$) at 42 hours. The metric used is the Jensen-Shannon Distance (JSD), as explained in the Supplementary Information (Section S4): a value of 00 indicates identical distributions, while $1$ indicates complete dissimilarity. Simulation setup: Setup2 in Section 3. The initial conditions are visualised in Figure S3 in Supplementary Information (Section S5), with $100$ cells, $\lambda_{1}+\lambda_{2}=0.5$, $a=1-\lambda_{2}/\lambda_{1}=0.9$, vertical $\hat{\mathbf{v}}_{1}$ and horizontal $\hat{\mathbf{v}}_{2}$.
@@
We also examine, given strong vertical contact guidance cues initially, how dominant the degradation phenotype must be for cells to effectively modulate the collagen fibre distribution and achieve isotropic invasion. To this end, we compare one-dimensional cell density profiles, generated by averaging the cell density in the $y$-direction, given different collagen degradation rates. The metric we use to compare these density profiles is the Jensen-Shannon distance (JSD), which is a bounded measure of how far apart two distributions are, with smaller values of the JSD indicating higher similarity (see Supplementary Information Section S4). Here, we take $\rm{JSD}=0.2$ as the threshold beyond which invasion is substantially different from the corresponding pattern observed with the same level of secretion but no degradation. Figure 5(c) shows that once the secretion rate exceeds a critical value ($s=2^{-7}$ $\rm{min}^{-1}$), the cell population cannot overcome the initial vertical contact guidance cues no matter how rapid the rate of degradation.
@@ 4.5 Showcase of invasion patterns across a range of different phenotypes
Finally, by varying specific parameter values listed in Table 2, we showcase the wide range of range of cell invasion and collagen fibre patterns that can arise due to different cell phenotypes. Figure 6(a) demonstrates that when cell motility is primarily driven by pairwise interactions rather than random motion, the colony remains compact with a high degree of radial symmetry. Because the cell motility is less stochastic in nature, the secreted collagen fibres exhibit a high degree of alignment. In contrast, when cell migration is dominated by random motility, the resulting cell cluster is more dispersed, and the secreted fibres appear less aligned. Another key observation is that finger-like invasion patterns emerge only when cells secrete collagen fibres sufficiently quickly (bottom row of Figure 6(a)). In this regime, cells at the edge of the cluster show a high degree of persistence, reinforcing both the orientation and anisotropy of the secreted fibres. This, in turn, further guides the direction of cell migration.
Figure 6(b) shows that varying the length of the averaging window (which determines the alignment of secreted collagen fibres) significantly influences the patterns of cell invasion. A longer averaging window smooths out the effects of random motion and results in the secretion of more aligned and spatially organised collagen fibres, which in turn enhances the emergence of a finger-like invasion pattern. On the other hand, a short averaging window leads to more random motility, and less well aligned collagen fibres. Additionally, the sensitivity of cells to contact guidance cues from the underlying collagen fibres plays a crucial role in shaping the patterns of cell invasion. If cells only respond to contact guidance cues when the collagen fibre area fraction is high (i.e. large $\bar{\lambda}$) then, despite an aligned and structured fibre field (due to the long averaging window), the cells do not exhibit finger-like behaviour.
@@ Acknowledgements
$\begin{split}\mathbf{\Omega}&=\lambda_{1}\hat{\mathbf{v}}_{1}\hat{\mathbf{v}}_{1}^{\mathrm{T}}+\lambda_{2}\hat{\mathbf{v}}_{2}\hat{\mathbf{v}}_{2}^{\mathrm{T}}\\ &=\begin{bmatrix}|&|\\ \hat{\mathbf{v}}_{1}&\hat{\mathbf{v}}_{2}\\ |&|\end{bmatrix}\begin{bmatrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{bmatrix}\begin{bmatrix}\mbox{---}&\hat{\mathbf{v}}_{1}&\mbox{---}\\ \mbox{---}&\hat{\mathbf{v}}_{2}&\mbox{---}\end{bmatrix}\\ &=\underbrace{\begin{bmatrix}\cos(\theta_{1})&-\sin(\theta_{1})\\ \sin(\theta_{1})&\cos(\theta_{1})\end{bmatrix}}_{:=\mathbf{P}\atop\text{anticlockwise rotation of }\theta_{1}}\underbrace{\begin{bmatrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{bmatrix}}_{:=\mathbf{D}\atop\text{rescale and re-orient}}\underbrace{\begin{bmatrix}\cos(\theta_{1})&\sin(\theta_{1})\\ -\sin(\theta_{1})&\cos(\theta_{1})\end{bmatrix}}_{:=\mathbf{P}^{\mathrm{T}}\atop\text{clockwise rotation of }\theta_{1}}.\end{split}$ (14)
@@
In all simulations, $100$ cells are initially arranged in a densely packed circular disc at the centre of the domain, forming a confluent configuration. For Figure 2 in the main text, we initialise the system without any collagen fibres. For Figure 3(a), (c), and (e), the collagen fibres are initialised with a high anisotropy degree ($a=0.9$), while the total fibre density ($\lambda_{1}+\lambda_{2}$) is varied among $0.1$, $0.4$, and $1$. For Figure 3(b), (d), and (f), the total fibre density is fixed at $\lambda_{1}+\lambda_{2}=0.8$, while the anisotropy degree is varied among $0.0$, $0.5$, and $0.95$. In Figure 4(a), collagen fibres are initialised with a high density ($\lambda_{1}+\lambda_{2}=0.9$), whereas in Figure 4(b), a low density ($\lambda_{1}+\lambda_{2}=0.1$) is used. In both cases, the fibres are aligned ($a=0.9$). For Figure 5, the collagen fibres are initialised with $\lambda_{1}+\lambda_{2}=0.5$ and $a=0.9$. In Figure 6, a sparse and isotropic collagen distribution is used, with $\lambda_{1}+\lambda_{2}=0.01$ and $a=0$. For Figures 25, the angle between the major fibre orientation and the positive $x$-axis, denoted $\angle\hat{\mathbf{v}}_{1}$, is sampled from the interval $\pi/2+U[-\pi/18,\pi/18]$. In contrast, for Figure 6, the collagen fibres are randomly oriented, with $\angle\hat{\mathbf{v}}_{1}$ sampled uniformly from $U[-\pi,\pi]$. All initial conditions, except for those in Figure 3, are visualised in Figure S3.
Figure S3: Initial conditions for results in the main text. Pale yellow backgrounds indicate regions devoid of collagen fibres. Simulation details: Setup1 with a larger domain and longer simulation time is used for Figure 2, while Setup2 is used for Figures 36.
S6 Collagen degrading phenotypes display different invasion patterns

### v2

@@ 4.2 Non-ECM-modulating phenotypes display restricted invasion into dense collagen environments
Another important prediction of the model is that the initial collagen fibre distribution indirectly influences cell proliferation. This occurs because the cell clusters are more compact when the collagen fibres are more dense and more aligned (Figure 3(c),(d), and Figure S2 in the Supplementary Information). This effect is not apparent until approximately $24$ hours from the start of the simulation (Figure 3(e),(f)), which corresponds to the average cell cycle length.
@@ 4.3 Collagen-modulating phenotypes display different invasion patterns
We now examine the dynamics of a collagen-secreting phenotype, setting the degradation rate to be zero, $d=0.0$, and the initial total area fraction of collagen to be low so that there is substantial area available into which cells can secrete new collagen. Here, the remodelling process is more intricate, and displays spatial variations across the cluster (see Figure 4(b)). For cells in the top and bottom regions (as defined in Figure 2), migration is primarily aligned with the vertically oriented major fibre direction. In these regions, anisotropy increases over time as newly secreted fibres reinforce this vertical alignment. For cells in the left and right regions (as defined in Figure 2), migration is primarily driven by population pressure arising from cell–cell interactions, and is aligned with the horizontally oriented minor fibre direction. As the cells secrete new collagen, which is primarily horizontally aligned, the fibre anisotropy initially decreases over time as the fibre distribution becomes more uniform (i.e., $\lambda_{1}\approx\lambda_{2}$). It then increases as newly secreted fibres are continually laid down in the horizontal alignment, so that the major fibre direction becomes horizontal. In the central region, the anisotropy decreases over time as cells move predominantly randomly. Therefore, as shown in Figure 4(b), the high collagen-secreting cell phenotype deposits denser collagen fibres in a more anisotropically aligned structure compared with the low-secreting phenotype. Moreover, high secretion rates result in clear major fibre orientations across the domain: horizontal in the left and right regions, and vertical in the top and bottom regions.
@@ 4.4 Tensioning secretion and degradation in the phenotype alters invasion patterns
Label Collagen-related cell phenotypes
$\star^{1}$ small secretion with large degradation rates
$\star^{2}$ large secretion and large degradation rates
$\star^{3}$ large secretion with small degradation rates
Table 4: Secretion and degradation cell phenotypes investigated in Section $4.$
@@
To investigate how a phenotype that both secretes and degrades fibres affects collagen fibre dynamics and cell invasiveness, we perform a parameter sweep across different fibre secretion and degradation rates. We consider a scenario in which there is very strong contact guidance initially due to a high area fraction of aligned collagen. Figure 5(a) illustrates the cell and collagen fibre distributions for small secretion with large degradation rates ($\star^{1}$), large secretion and large degradation rates ($\star^{2}$), and large secretion with small degradation rates ($\star^{3}$), see Table 4. For a cell phenotype that predominantly degrades collagen ($\star^{1}$), the initial collagen distribution is degraded and becomes isotropically distributed, reducing the strength of contact guidance while facilitating cell invasion in all directions. This is consistent with results in Figure 4(a). For a cell phenotype that equally secretes and degrades collagen ($\star^{2}$), the initially aligned fibres are degraded: this results in reduced contact guidance so that newly secreted collagen is relatively isotropic. Comparing the $\star^{2}$ phenotype with $\star^{1}$, we observe that migration is facilitated in all directions, with a stronger tendency in the vertical direction. This arises because the higher secretion rate of the $\star^{2}$ phenotype produces denser collagen fibres than the $\star^{1}$ phenotype. However, the large degradation rate of the $\star^{2}$ phenotype subsequently leads to a weakening of these vertical cues, thereby eventually facilitating horizontal migration. Finally, for a cell phenotype that predominantly secretes collagen ($\star^{3}$), the initial anisotropy, and thus the contact guidance cues, are reinforced in the top and bottom regions. This occurs because cells in these regions tend to migrate vertically due to population pressure, even in the absence of contact guidance. This promotes enhanced invasion along the vertical axis, while horizontal invasion is suppressed because guidance cues from population pressure are in tension with those from the collagen distribution. This is consistent with results in Figure 4(b).
@@
Figure 5: Tensioning secretion and degradation in the phenotype alters invasion patterns. Results from a parameter sweep across secretion ($s$) and degradation ($d$) rates. (a) Cell and collagen fibre distributions at $42$ hours with $\star^{1}$$\left(s,d\right)=\left(2^{-10},2^{-4}\right)$ $\rm{min}^{-1}$, $\star^{2}$$\left(s,d\right)=\left(2^{-4},2^{-4}\right)$ $\rm{min}^{-1}$, and $\star^{3}$$\left(s,d\right)=\left(2^{-4},2^{-10}\right)$ $\rm{min}^{-1}$, respectively. (b) Heat map showing the similarity of the average cell density along the horizontal direction, compared to that of the non-degrading cell phenotype ($d=0,\rm{min}^{-1}$) at 42 hours. The metric used is the Jensen-Shannon Distance (JSD), as explained in the Supplementary Information (Section S4): a value of 0 indicates identical distributions, while 11 indicates complete dissimilarity. Simulation setup: Setup2 in Section 3. The initial conditions are visualised in Figure S3 in Supplementary Information (Section S5), with $100$ cells, $\lambda_{1}+\lambda_{2}=0.5$, $a=1-\lambda_{2}/\lambda_{1}=0.9$, vertical $\hat{\mathbf{v}}_{1}$ and horizontal $\hat{\mathbf{v}}_{2}$.
@@
We also examine, given strong vertical contact guidance cues initially, how dominant the degradation phenotype must be for cells to effectively modulate the collagen fibre distribution and achieve isotropic invasion. To this end, we compare one-dimensional cell density profiles, generated by averaging the cell density in the $y$-direction, given different collagen degradation rates. The metric we use to compare these density profiles is the Jensen-Shannon distance (JSD), which is a bounded measure of how far apart two distributions are, with smaller values of the JSD indicating higher similarity (see Supplementary Information Section S4). Here, we take $\rm{JSD}=0.2$ as the threshold beyond which invasion is substantially different from the corresponding pattern observed with the same level of secretion but no degradation. Figure 5(b) shows that once the secretion rate exceeds a critical value ($s=2^{-7}$ $\rm{min}^{-1}$), the cell population cannot overcome the initial vertical contact guidance cues no matter how rapid the rate of degradation.
@@ 4.5 Showcase of invasion patterns across a range of different phenotypes
Finally, by varying specific parameter values listed in Table 2, we showcase the wide range of cell invasion and collagen fibre patterns that can arise due to different cell phenotypes. Figure 6(a) demonstrates that when cell motility is primarily driven by pairwise interactions rather than random motion, the colony remains compact with a high degree of radial symmetry. Because the cell motility is less stochastic in nature, the secreted collagen fibres exhibit a high degree of alignment. In contrast, when cell migration is dominated by random motility, the resulting cell cluster is more dispersed, and the secreted fibres appear less aligned. Another key observation is that finger-like invasion patterns emerge only when cells secrete collagen fibres sufficiently quickly (bottom row of Figure 6(a)). In this regime, cells at the edge of the cluster show a high degree of persistence, reinforcing both the orientation and anisotropy of the secreted fibres. This, in turn, further guides the direction of cell migration.
Figure 6(b) shows that varying the length of the averaging window (which can be referred to as the memory window, and determines the alignment of secreted collagen fibres) significantly influences the patterns of cell invasion. A longer averaging window (i.e. longer memory) smooths out the effects of random motion and results in the secretion of more aligned and spatially organised collagen fibres, which in turn enhances the emergence of a finger-like invasion pattern. On the other hand, a short averaging window (i.e. short memory) leads to more random motility, and less well aligned collagen fibres. Additionally, the sensitivity of cells to contact guidance cues from the underlying collagen fibres plays a crucial role in shaping the patterns of cell invasion. If cells only respond to contact guidance cues when the collagen fibre area fraction is high (i.e. large $\bar{\lambda}$) then, despite an aligned and structured fibre field (due to the long averaging window), the cells do not exhibit finger-like behaviour.
@@ Acknowledgements
[…]
@@
[…]
S6 Different collagen modulating kernels $\omega$ lead to similar dynamics
Recall that the collagen-modulating kernel $\omega$, defined in Equation (2), characterises the strength and range of collagen degradation and secretion. We now investigate whether the cell and collagen fibre dynamics are sensitive to the functional form of $\omega$. Figure S4 illustrates cell migration and collagen fibre dynamics when $\omega$ is taken to be a Heaviside function:
$\omega\left(\mathbf{X}^{i},\mathbf{x},\sigma\right)=\begin{cases}1,&\quad|\mathbf{X}^{i}-\mathbf{x}|\leq\sigma/2,\\ 0,&\quad|\mathbf{X}^{i}-\mathbf{x}|>\sigma/2.\end{cases}$ (18)
We note that since $\sigma/2$ is a measure of the cell radius, this functional form is intended to restrict secretion and degradation locally to a cell. On the other hand, Figure S5 shows the cell migration and collagen fibre dynamics for
$\omega\left(\mathbf{X}^{i},\mathbf{x},\sigma\right)=\begin{cases}1-\frac{|\mathbf{X}^{i}-\mathbf{x}|}{\sigma},&\quad|\mathbf{X}^{i}-\mathbf{x}|\leq\sigma,\\ 0,&\quad|\mathbf{X}^{i}-\mathbf{x}|>\sigma.\end{cases}$ (19)
In this case the range over which collagen is secreted and degraded is doubled to $\sigma$. Comparing Figures S4S5 with Figure 2, we observe no qualitative differences in cell or collagen fibre dynamics.
Figure S4: Cell migration patterns are shown when the collagen-modulating kernel $\omega$ is taken as a Heaviside function with support $\sigma/2$, corresponding to the cell radius. This figure can be compared with Figure 2 in the main text.
Figure S5: Cell migration patterns are shown when the range of collagen secretion and degradation extends beyond the space occupied by a single cell; that is, the range of collagen modulation is $\sigma$, with $\sigma/2$ representing the cell radius. This figure can be compared with Figure 2 in the main text.
S7 Collagen degrading phenotypes display different invasion patterns